5/y+2 minus 3/y-2 =4/y^2-4
Answer by feliz1965(151)

The LCD is the denominator on the right side of the equation.
Multiply each term by the LCD as step one.
5/(y + 2) times (y^2 - 4) = 5(y - 2)
3/(y - 2) times (y^2 - 4) = -3(y + 2)
4/(y^2 - 4) times (y^2 - 4) = 4
We now have a linear equation:
5(y - 2) -3(y + 2) = 4
Can you finish now?